Find the common factors of the given terms.
10 pq, 20qr, 30 rp

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 10 pq, 20qr and 30rp
10 pq = 2 × 5 × p × q
20qr = 2 × 2 × 5 × q × r
30rp= 2 × 3 × 5 × r × p
Common factors of 10 pq, 20qr and 30rp are 2, 5 and, 2×5 = 10

Thanks